A police investigator measures straight skid marks 27 m long in an accident investigation. Assuming a friction force is 8.00 x10^3 and a car mass is 1.00 x10^3 kg, what was the minimum speed of the car when the brakes locked?


1
Expert's answer
2022-05-20T13:19:31-0400

The energy-work theorem says

"\frac{mv^2}{2}=Fd"

Thus

"v=\sqrt{2Fd/m}\\
=\sqrt{2*8*10^3*27/1.0*10^3}=20.8\:\rm m/s"